OpenStudy (anonymous):

Help Please!! A circle is inscribed in a regular hexagon with each side 8.0 meters long. Find the radius of the circle.

OpenStudy (dumbcow):

 from the right triangle tan(60) = r/4 sqrt3 = r/4 --> r = 4sqrt3

OpenStudy (anonymous):

wow, thanks alot @dumbcow . :D